# Guiding diffusion models to reconstruct flow fields from sparse data

### v1

@@ Abstract
The reconstruction of unsteady flow fields from limited measurements is a challenging and crucial task for many engineering applications. Machine learning models are gaining popularity in solving this problem due to their ability to learn complex patterns from data and generalize across diverse conditions. Among these, diffusion models have emerged as particularly powerful in generative tasks, producing high-quality samples by iteratively refining noisy inputs. In contrast to other methods, these generative models are capable of reconstructing the smallest scales of the fluid spectrum. In this work, we introduce a novel sampling method for diffusion models that enables the reconstruction of high-fidelity samples by guiding the reverse process using the available sparse data. Moreover, we enhance the reconstructions with available physics knowledge using a conflict-free update method during training. To evaluate the effectiveness of our method, we conduct experiments on 2 and 3-dimensional turbulent flow data. Our method consistently outperforms other diffusion-based methods in predicting the fluid’s structure and in pixel-wise accuracy. This study underscores the remarkable potential of diffusion models in reconstructing flow field data, paving the way for their application in Computational Fluid Dynamics research.
@@ I Introduction
Fluid dynamics plays an essential role in many modern engineering applications, including vehicle and turbine design Pfeiffer (2016); Strom, Brunton, and Polagye (2017); Maurya, Maurya, and Luby (2018), active flow control Manohar et al. (2018); Brunton and Noack (2015); Sipp and Schmid (2016), cardiac blood flow modelling Yakhot, Anor, and Karniadakis (2007); Sankaran et al. (2012), ship wake detection Graziano, D’Errico, and Rufino (2016) and climate predictions Kalnay (2003). Numerical simulations and experimental studies are key methods of investigating complex flow phenomena. However, both approaches have inherent limitations. Although advancements in computational fluid dynamics (CFD) and hardware capabilities now enable high-fidelity simulations PI et al. (2013), accurate simulation of large-scale industrial problems still requires significant computational resources Carlberg et al. (2019). Thus, reconstructing fine flow structures from sparse data is still a critical topic for fluid simulations. Meanwhile, experimental studies are often limited by sensor numbers in the flow field, and evaluating complex flow behaviour through sparse signals from sensors is a huge challenge for experimental fluid dynamics Qin et al. (2022); Raffel et al. (2018).
Sparse reconstruction is a technique employed to recover or approximate the full-scale features of a system using only a limited subset of data, such as selected measurements, pixels, or observations. It has been pivotal in aerodynamic applications, as seen in early work of Bui-Thanh et al. Bui-Thanh, Damodaran, and Willcox (2004), where they utilized Proper Orthogonal Decomposition (POD) to resolve complex flow fields and aerodynamic forces efficiently. Modifications of this method like Gappy-POD Venturi and Karniadakis (2004); Raben, Charonko, and Vlachos (2012) or other linear methods such as Dynamic Mode Decomposition (DMD) Schmid (2010); Tu (2013) have been used to extract the coherent structures and correlation features of a fluid flow, which can later be used to reconstruct its high-resolution representation. A different approach was used in the work of Callaham et al. Callaham, Maeda, and Brunton (2018), where sparse representation was leveraged to reconstruct target states from a dictionary of precomputed examples, showing better robustness than the previously introduced methods when encountering noisy data. Loiseau et al. Loiseau, Noack, and Brunton (2018) presented a sparse reduced-order modeling framework that combined feature extraction, sparse nonlinear dynamics identification, and modal decomposition. They used it to reconstruct full-state fluid dynamics from typical experimental data such as time-resolved sensor data and optionally also non-time-resolved Particle Image Velocimetry (PIV) snapshots. The article from Zaki et al. Zaki and Wang (2021) proposed a hybrid approach bridging physics-based modelling and data-driven optimization, where an adjoint-variational method is introduced to reconstruct turbulent flow.
@@
Solving the sparse reconstruction problem can also be seen as a super-resolution task, where our set of sparse sensor measurements compose a low-resolution sample, and we aim to reconstruct its high-resolution counterpart Fukami, Fukagata, and Taira (2023). Sun et al. Sun and Wang (2020) proposed a physics-constrained Bayesian neural network (PC-BNN) to reconstruct a flow field, modelling their problem as a super-resolution task. Fukami et al. Fukami et al. (2021) took this same approach by computing a Voronoi tessellation of the sparse information and then using it as an input for a CNN, an established architecture for image super-resolution Dong et al. (2015). Another example is the work of Güemes et al. Güemes, Sanmiguel Vila, and Discetti (2022), where they presented a framework called RaSeedGAN that interpolated the scattered data onto a regular grid and afterwards upsampled it using a Generative Adversarial Network (GAN).
@@
In recent years, diffusion models Sohl-Dickstein et al. (2015); Song et al. (2020); Ho, Jain, and Abbeel (2020) have shown superior performance in tasks like image generation Dhariwal and Nichol (2021), image editing Meng et al. (2021); Avrahami, Lischinski, and Fried (2022), image inpainting Lugmayr et al. (2022); Chung, Sim, and Ye (2022), and image colorization Saharia et al. (2022a); Yue, Wang, and Loy (2024). Following this success, there has been a rising interest in solving image super-resolution tasks with these models to improve in reconstruction accuracy. A direct approach to achieve this goal is to give the low-resolution image as an input to the model, and train it to map the high-resolution counterpart Saharia et al. (2022b); Rombach et al. (2022). Other approaches utilize a pre-trained diffusion model and only modify the sampling methods Chung, Sim, and Ye (2022); Choi et al. (2021); Yue and Loy (2024); Wang et al. (2024). More recently, diffusion models have also been used to predict physical flows. Liu et al. Liu and Thuerey (2024) explored their use as uncertainty-aware surrogate models for turbulence simulations, demonstrating their ability to capture full solution distributions and outperform traditional probabilistic approaches. Moreover, Shu et al. Shu, Li, and Farimani (2023) presented a super-resolution method for reconstructing high-fidelity flow fields using this same architecture. Diffusion models have also been used to predict nonlinear fluid fields from discrete simulation data Yang and Sommer (2023), turbulent rotating flows Li et al. (2023), and supersonic flows Abaidi and Adams (2025).
With their generative capabilities and the proficiency to accurately learn complex distributions, diffusion models emerged as a promising candidate to enhance the existing reconstruction methods found in the literature. They are especially interesting in combination with physical priors, such as conservation laws.
@@
In terms of model architectures, early approaches primarily relied on CNNs applied to cartesian grids Thuerey et al. (2020). Later, neural operators emerged, most notably in the form of Fourier Neural Operators (FNO) Li et al. (2021), and Graph Neural Networks (GNNs) Battaglia et al. (2016); Lino et al. (2022). More recently, spatial transformer architectures were employed as neural operators, e.g., via OFormer Li, Meidani, and Farimani (2022), Universal Physics Transformers (UPT) Alkin et al. (2025) and Transolver Wu et al. (2024), among others. Transformers have experienced particular interest due to their ability to capture global and multi-scale interactions. Their main drawback is the computational $N^{2}$ scaling of their inherent attention mechanism, where N is the number of grid nodes. Thus, vanilla attention can be prohibitive for large 2D or 3D problems. The P3D model Holzschuh et al. (2025), which we employ in our experiments, employs shifted-window attention Liu et al. (2021) to improve on this behavior. Even with architectural advances, the computational cost of handling high-resolution 3D fields has restricted applications to low-resolution 3D settings with limited accuracy. This includes deterministic models Zhou et al. (2022); Maurya (2025), GANs Subramaniam et al. (2020); Kim et al. (2021), and diffusion approaches Bi et al. (2025); Huang et al. (2024). In this context, our work advances diffusion-based super-resolution methods to 3D fields with high resolutions.
@@ II.1 Diffusion models
In this work, we aim to reconstruct a high-fidelity sample $\mathbf{x_{0}}\in\mathbb{R}^{n\times n}$ from a subset of sparse true values $\mathbf{x_{s}}$ with denoising diffusion probabilistic models (DDPMs) Sohl-Dickstein et al. (2015); Ho, Jain, and Abbeel (2020). DDPMs are generative models that learn to gradually transform samples from a simple known distribution $p(\mathbf{x}_{T})$ into samples from the target distribution $p(\mathbf{x}_{0})$ that we want to model. Typically, $p(\mathbf{x}_{T})$ is a normal distribution $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The transformation from $p(\mathbf{x}_{T})$ to $p(\mathbf{x}_{0})$, usually known as reverse process, is an iterative process described by the learnable distribution $p_{\theta}(\mathbf{x}_{0:T})$:
@@
With the above forward procedure, Ho et al. Ho, Jain, and Abbeel (2020) proposed a simple loss function where the network directly learns the noise $\epsilon$ in the forward procedure:
@@ II.3 Physics-based Diffusion Training
where $\lambda$ is a balance factor for the physical residual. Similar methods have been proposed by previous works Bastek, Sun, and Kochmann (2024); Shan et al. (2024), such as the Physics-Informed Diffusion Models (PIDMs) Bastek, Sun, and Kochmann (2024). In PIDM, a central difficulty is that the scaling factor is highly dependent on the specificity of the training task at hand. We consider an alternative where the rescaling factor is dynamically set to balance the importance of each term of the global loss. This balance is computed via the ratio of the noise step and the physics residual loss as:
@@
To improve on these manually-scaled and heuristic approaches for balancing, we build on a recent method for conflict-free updates, ConFIG Liu, Chu, and Thuerey (2024), to automatically balance the loss terms. Defining $\boldsymbol{g}_{D}={\partial\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|_{2}}/{\partial\theta}$ and $\boldsymbol{g}_{F}={\partial\|\mathcal{F}(\tilde{\mathbf{x}}_{0}^{(t)})\|_{2}}/{\partial\theta}$ as the gradients for diffusion loss and physical residual loss terms, the ConFIG method returns a conflict-free update direction $\boldsymbol{g}_{c}$ as
@@
Provably, $\boldsymbol{g}_{D}^{\top}\boldsymbol{g}_{c}>0$ and $\boldsymbol{g}_{F}^{\top}\boldsymbol{g}_{c}>0$ hold, meaning that optimization along $\boldsymbol{g}_{c}$ will decrease both the diffusion loss and physical residual loss simultaneously. The generalization of the operator to more than two gradients can be found in Liu’s et al. articleLiu, Chu, and Thuerey (2024). As shown in equation (15), the magnitude of $\boldsymbol{g}_{v}$ in the ConFIG method is rescaled based on the sum of each loss-specific gradient’s projection on the conflict-free direction. However, for sparse reconstruction tasks, the physical residual loss is usually significantly larger than the diffusion loss. This leads to the magnitude change of the $\boldsymbol{g}_{v}$ mainly tracking the magnitude change of the residual gradient. For momentum-based optimizers like Adam, this represents a problem, as these optimizers adaptively change the learning rate according to the variation of the gradient. In turn, it makes the balancing approach of the ConFIG method ineffective. To counteract this undesirable behavior, we propose the following modification to the ConFIG method, which computes the magnitude of the resulting gradient after normalizing both loss-specific gradients:
@@
This results in a method that considers both loss terms equally for the computation of the gradient magnitude. At the same time, it also allows the Adam optimizer to compute the moments of the gradient in a balanced manner. The caveat of this modification is that we do not transfer any information about the original magnitude of $\boldsymbol{g}_{D}$ and $\boldsymbol{g}_{F}$ to the optimizer, as the new magnitude only contains the information of how similar the direction of $\boldsymbol{g}_{v}$ is to the directions of both loss terms. Nonetheless, our results show that the benefits of conflict-free learning updates clearly outweigh the lack of information about the original magnitudes.
We call the resulting method $\text{ConFIG}_{u}$ in the following, and Figure 3 illustrates how our update strategy differs from the original ConFIG method for a specific two-dimensional example. The $\text{ConFIG}_{\text{u}}$ updates in blue preserve significantly more information from the diffusion updates and are generally more balanced. They effectively capture the agreement between the directions of the two loss gradients. We compare the performance of these training strategies in Section III.2.
@@ III Results and discussion
In this section, we evaluate our method in two challenging scenarios: 2D Kolmogorov flow and 3D isotropic turbulence. We first compare the performance of different training methods, demonstrating how our $\text{ConFIG}_{\text{u}}$ training method outperforms existing approaches in generating physically consistent samples. Next, we investigate the effectiveness of our proposed masked diffusion sampling method for sparse data reconstruction. The results show that our method effectively reconstructs fine-grained details and preserves the structural integrity of the flow fields, outperforming baseline methods, especially in capturing high-frequency components that are critical in turbulent flows. Finally, we provide a qualitative and quantitative analysis of the reconstructed samples, including visualizations and an evaluation of the vorticity and Q-criterion distributions.

### v2

@@ Abstract
The reconstruction of unsteady flow fields from limited measurements is a challenging and crucial task for many engineering applications. Machine learning models are gaining popularity for solving this problem due to their ability to learn complex patterns from data and to generalize across diverse conditions. Among these, diffusion models have emerged as being particularly powerful for generative tasks, producing high-quality samples by iteratively refining noisy inputs. In contrast to other methods, these generative models are capable of reconstructing the smallest scales of the fluid spectrum. In this work, we introduce a novel sampling method for diffusion models that enables the reconstruction of high-fidelity samples by guiding the reverse process using the available sparse data. Moreover, we enhance the reconstructions with available physics knowledge using a conflict-free update method during training. To evaluate the effectiveness of our method, we conduct experiments on 2 and 3-dimensional turbulent flow data. Our method consistently outperforms other diffusion-based methods in predicting the fluid’s structure and in pixel-wise accuracy. This study underscores the remarkable potential of diffusion models in reconstructing flow field data, paving the way for leveraging them in fluid dynamics research and applications ranging from super-resolution to reconstructions of experiments.
@@ I Introduction
Fluid dynamics plays an essential role in many modern engineering applications, including vehicle and turbine design Pfeiffer (2016); Strom et al. (2017); Maurya et al. (2018), active flow control Manohar et al. (2018); Brunton and Noack (2015); Sipp and Schmid (2016), cardiac blood flow modelling Yakhot et al. (2007); Sankaran et al. (2012), ship wake detection Graziano et al. (2016) and climate predictions Kalnay (2003). Numerical simulations and experimental studies are key methods of investigating complex flow phenomena. However, both approaches have inherent limitations. Although advancements in computational fluid dynamics (CFD) and hardware capabilities now enable high-fidelity simulations PI et al. (2013), accurate simulation of large-scale industrial problems still requires significant computational resources Carlberg et al. (2019). Thus, reconstructing fine flow structures from sparse data is still a critical topic for fluid simulations. Meanwhile, experimental studies are often limited by sensor numbers in the flow field, and evaluating complex flow behaviour through sparse signals from sensors is a huge challenge for experimental fluid dynamics Qin et al. (2022); Raffel et al. (2018).
Sparse reconstruction is a technique employed to recover or approximate the full-scale features of a system using only a limited subset of data, such as selected measurements, pixels, or observations. It has been pivotal in aerodynamic applications, as seen in early work of Bui-Thanh et al. Bui-Thanh et al. (2004), where they utilized Proper Orthogonal Decomposition (POD) to resolve complex flow fields and aerodynamic forces efficiently. Modifications of this method like Gappy-POD Venturi and Karniadakis (2004); Raben et al. (2012) or other linear methods such as Dynamic Mode Decomposition (DMD) Schmid (2010); Tu (2013) have been used to extract the coherent structures and correlation features of a fluid flow, which can later be used to reconstruct its high-resolution representation. A different approach was used in the work of Callaham et al. Callaham et al. (2018), where sparse representation was leveraged to reconstruct target states from a dictionary of precomputed examples, showing better robustness than the previously introduced methods when encountering noisy data. Loiseau et al. Loiseau et al. (2018) presented a sparse reduced-order modeling framework that combined feature extraction, sparse nonlinear dynamics identification, and modal decomposition. They used it to reconstruct full-state fluid dynamics from typical experimental data such as time-resolved sensor data and optionally also non-time-resolved Particle Image Velocimetry (PIV) snapshots. The article from Zaki et al. Zaki and Wang (2021) proposed a hybrid approach bridging physics-based modelling and data-driven optimization, where an adjoint-variational method is introduced to reconstruct turbulent flow.
@@
Solving the sparse reconstruction problem can also be seen as a super-resolution task, where our set of sparse sensor measurements compose a low-resolution sample, and we aim to reconstruct its high-resolution counterpart Fukami et al. (2023). Sun et al. Sun and Wang (2020) proposed a physics-constrained Bayesian neural network (PC-BNN) to reconstruct a flow field, modelling their problem as a super-resolution task. Fukami et al. Fukami et al. (2021) took this same approach by computing a Voronoi tessellation of the sparse information and then using it as an input for a CNN, an established architecture for image super-resolution Dong et al. (2015). Another example is the work of Güemes et al. Güemes et al. (2022), where they presented a framework called RaSeedGAN that interpolated the scattered data onto a regular grid and afterwards upsampled it using a Generative Adversarial Network (GAN).
@@
In recent years, diffusion models Sohl-Dickstein et al. (2015); Song et al. (2020); Ho et al. (2020) have shown superior performance in tasks like image generation Dhariwal and Nichol (2021), image editing Meng et al. (2021); Avrahami et al. (2022), image inpainting Lugmayr et al. (2022); Chung et al. (2022b), and image colorization Saharia et al. (2022a); Yue et al. (2024). Following this success, there has been a rising interest in solving image super-resolution tasks with these models to improve in reconstruction accuracy. A direct approach to achieve this goal is to give the low-resolution image as an input to the model, and train it to map the high-resolution counterpart Saharia et al. (2022b); Rombach et al. (2022). Other approaches utilize a pre-trained diffusion model and only modify the sampling methods Chung et al. (2022b); Choi et al. (2021); Yue and Loy (2024); Wang et al. (2024). More recently, diffusion models have also been used to predict physical flows. Liu et al. Liu and Thuerey (2024) explored their use as uncertainty-aware surrogate models for turbulence simulations, demonstrating their ability to capture full solution distributions and outperform traditional probabilistic approaches. Moreover, Shu et al. Shu et al. (2023) presented a super-resolution method for reconstructing high-fidelity flow fields using this same architecture. Diffusion models have also been used to predict nonlinear fluid fields from discrete simulation data Yang and Sommer (2023), turbulent rotating flows Li et al. (2023), and supersonic flows Abaidi and Adams (2025). With their generative capabilities and the proficiency to accurately learn complex distributions, diffusion models emerged as a promising candidate to enhance the existing reconstruction methods found in the literature. They are especially interesting in combination with physical priors, such as conservation laws. These methods can also be adapted to solve inverse problems, similar to our reconstruction problem. One example is Diffusion Posterior Sampling (DPS) Chung et al. (2022a), which modifies the score function of diffusion models to make the sample closer to the observation during inference. This can lead to better accuracy, with the caveat of higher computational cost at inference time, as backpropagation through the network has to be computed in every step of the generation process. Furthermore, DPS only enforces the observed values as a soft constraint by steering the probability flow according to the observations. In our algorithm, we seek to address these limitations, avoiding backpropagation during the generation and providing a stronger constraint over the observations.
@@
In terms of model architectures, early approaches primarily relied on CNNs applied to Cartesian grids Thuerey et al. (2020). Later, neural operators emerged, most notably in the form of Fourier Neural Operators (FNO) Li et al. (2021), and Graph Neural Networks (GNNs) Battaglia et al. (2016); Lino et al. (2022). More recently, spatial transformer architectures were employed as neural operators, e.g., via OFormer Li et al. (2022), Universal Physics Transformers (UPT) Alkin et al. (2025) and Transolver Wu et al. (2024), among others. Transformers have experienced particular interest due to their ability to capture global and multi-scale interactions. Their main drawback is the computational $N^{2}$ scaling of their inherent attention mechanism, where N is the number of grid nodes. Thus, vanilla attention can be prohibitive for large 2D or 3D problems. The P3D model Holzschuh et al. (2025), which we employ in our experiments, employs shifted-window attention Liu et al. (2021) to improve on this behavior. Even with architectural advances, the computational cost of handling high-resolution 3D fields has restricted applications to low-resolution 3D settings with limited accuracy. This includes deterministic models Zhou et al. (2022); Maurya (2025), GANs Subramaniam et al. (2020); Kim et al. (2021), and diffusion approaches Bi et al. (2025); Huang et al. (2024). In this context, our work advances diffusion-based super-resolution methods to 3D fields with high resolutions.
@@ II.1 Diffusion models
In this work, we aim to reconstruct a high-fidelity sample $\mathbf{x_{0}}\in\mathbb{R}^{n\times n}$ from a subset of sparse true values $\mathbf{x_{s}}$ with denoising diffusion probabilistic models (DDPMs) Sohl-Dickstein et al. (2015); Ho et al. (2020). DDPMs are generative models that learn to gradually transform samples from a simple known distribution $p(\mathbf{x}_{T})$ into samples from the target distribution $p(\mathbf{x}_{0})$ that we want to model. Typically, $p(\mathbf{x}_{T})$ is a normal distribution $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The transformation from $p(\mathbf{x}_{T})$ to $p(\mathbf{x}_{0})$, usually known as reverse process, is an iterative process described by the learnable distribution $p_{\theta}(\mathbf{x}_{0:T})$:
@@
With the above forward procedure, Ho et al. Ho et al. (2020) proposed a simple loss function where the network directly learns the noise $\epsilon$ in the forward procedure:
@@ II.3 Physics-based Diffusion Training
where $\lambda$ is a balance factor for the physical residual. Similar methods have been proposed by previous works Bastek et al. (2024); Shan et al. (2024), such as the Physics-Informed Diffusion Models (PIDMs) Bastek et al. (2024). In PIDM, a central difficulty is that the scaling factor is highly dependent on the specificity of the training task at hand. We consider an alternative approach where the rescaling factor is dynamically adjusted to balance the importance of each term in the global loss. This balance is computed via the ratio of the noise step and the physics residual loss as:
@@
To improve on these manually-scaled and heuristic approaches for balancing, we build on a recent method for conflict-free updates, ConFIG Liu et al. (2024), to automatically balance the loss terms. Defining $\boldsymbol{g}_{D}={\partial\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|_{2}}/{\partial\theta}$ and $\boldsymbol{g}_{F}={\partial\|\mathcal{F}(\tilde{\mathbf{x}}_{0}^{(t)})\|_{2}}/{\partial\theta}$ as the gradients for diffusion loss and physical residual loss terms, the ConFIG method returns a conflict-free update direction $\boldsymbol{g}_{c}$ as
@@
Provably, $\boldsymbol{g}_{D}^{\top}\boldsymbol{g}_{c}>0$ and $\boldsymbol{g}_{F}^{\top}\boldsymbol{g}_{c}>0$ hold, meaning that optimization along $\boldsymbol{g}_{c}$ will decrease both the diffusion loss and physical residual loss simultaneously. The generalization of the operator to more than two gradients is presented in Liu et al.’s article Liu et al. (2024). As shown in equation (15), the magnitude of $\boldsymbol{g}_{v}$ in the ConFIG method is rescaled based on the sum of each loss-specific gradient’s projection on the conflict-free direction. However, for sparse reconstruction tasks, the physical residual loss is usually significantly larger than the diffusion loss. This leads to the magnitude change of the $\boldsymbol{g}_{v}$ mainly tracking the magnitude change of the residual gradient. For momentum-based optimizers like Adam, this represents a problem, as these optimizers adaptively change the learning rate according to the variation of the gradient. In turn, it makes the balancing approach of the ConFIG method ineffective. To counteract this undesirable behavior, we propose the following modification to the ConFIG method, which computes the magnitude of the resulting gradient after normalizing both loss-specific gradients:
@@
This results in a method that considers both loss terms equally when computing the gradient magnitude. At the same time, it also allows the Adam optimizer to compute the moments of the gradient in a balanced manner. The caveat of this modification is that we do not transfer any information about the original magnitude of $\boldsymbol{g}_{D}$ and $\boldsymbol{g}_{F}$ to the optimizer, as the new magnitude only contains the information of how similar the direction of $\boldsymbol{g}_{v}$ is to the directions of both loss terms. Nonetheless, our results show that the benefits of conflict-free learning updates clearly outweigh the lack of information about the original magnitudes.
We refer to the resulting method as $\text{ConFIG}_{u}$ in the following, and Figure 3 illustrates how our update strategy differs from the original ConFIG method for a specific two-dimensional example. The $\text{ConFIG}_{\text{u}}$ updates in blue preserve significantly more information from the diffusion updates and are generally more balanced. They effectively capture the agreement between the directions of the two loss gradients. We compare the performance of these training strategies in Section III.2.
@@ III Results and discussion
In this section, we evaluate our method in two challenging scenarios: 2D Kolmogorov flow and 3D isotropic turbulence. We first compare the performance of different training methods, demonstrating how our $\text{ConFIG}_{\text{u}}$ training method outperforms existing approaches in generating physically consistent samples. Next, we investigate the effectiveness of our proposed masked diffusion sampling method for sparse data reconstruction. The results demonstrate how our method effectively reconstructs fine-grained details and preserves the structural integrity of the flow fields, outperforming baseline methods, particularly in capturing high-frequency components that are crucial in turbulent flows. Finally, we provide a qualitative and quantitative analysis of the reconstructed samples, including visualizations and an evaluation of the vorticity and Q-criterion distributions.